OpenStudy (anonymous):

h(x)=(g o f)(x)=1/(x+3)^2 which of the fallowing could be a possible decomposition of h(x)? A. f(x)=1/x^2;g(x)=x+3 B. f(x)=x+3; g(x)=1/x^2 C.f(x)=x^2; g(x)=x+3 D. f(x)=1/x;g(x)=x+3

OpenStudy (anonymous):

B

OpenStudy (anonymous):

coulld you explain how you got that so i could understand for future reference

OpenStudy (anonymous):

of course =) when you have (g o f)(x) it means do f(x) first and plug that answer into g(x)

OpenStudy (anonymous):

if you take B f(x)=x+3 g(x)=1/x^2 plug the f(x) into the g(x) formula (in other words, f(x) becomes the x for g) g(x)=1/x^2 g(x)=1/(x+3)^2

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

anytime =)